x varies directly as the square of s and inversely as t. how does x change when s is doubled? When both s and t are doubled?
Answer by stanbon(75887)

x varies directly as the square of s and inversely as t. how does x change when s is doubled? When both s and t are doubled?
------------------
Initial Conditions:
x = ks^2/t
---------------
When s is doubled:
x = k(2s)^2/t
x = k(4s^2)/t
x = 4[ks^2/t]
Conclusion: x is multiplied by 4
----------------------
When both s and t are doubled:
x = k(2s)^2/(2t)
x = k*4s^2/2t
x = k*2s^2/t
x = 2[ks^2/t]
-------
Conclusion: x is doubled.
================
Cheers,
Stan H.